Maintenance costs at Seaside Manufacturing over the past six months are listed in the following table. Month Maintenance cost Machine hours January ​$13,600 ​15,500 February ​$14,720 ​16,900 March $ 12 comma 000 13 comma 000 April ​$14,480 ​16,600 May $ 15 comma 000 17 comma 000 June ​$13,200 ​15,000 Using the high minus low ​method, what would the monthly fixed maintenance cost​ be? (Round any intermediary calculations to the nearest​ cent.)
Answer #1

High low Method

1.) Variable cost per unit = [Highest level cost - Lowest level cost] / [ Highest level hours - Lowest level hour]

=[$15000 - $12000] / [17000 units - 13000 units]

= $3000 / 4000

= $0.75 per hour

2.)      Fixed cost per month

Highest level cost - variable cost * Highest level hour = Lowest level cost - variable cost *Lowest level hour

$15000 - ($0.75 * 17000) = $12000 - ($0.75 * 13000 units)

$2250 = 2250

fixed maintenance cost per month = $2250
Answer #2

Answer:

Monthly Fixed Maintenance Cost = $4,400


Steps (High-Low Method):

  1. Identify High/Low Activity:

    • High: May (17,000 machine hours, $15,000 cost).

    • Low: March (13,000 machine hours, $12,000 cost).

  2. Calculate Variable Cost per Unit:

    Variable Cost=ΔCostΔHours=15,00012,00017,00013,000=3,0004,000=$0.75per machine hour

  3. Solve for Fixed Cost (Using High Point):

    Fixed Cost=Total Cost(Variable Cost×Hours)=15,000(0.75×17,000)=15,00012,750=$2,250

     

    Correction: Using the correct high point (June vs. March):

    Variable Cost=13,20012,00015,00013,000=$0.60per hourFixed Cost=13,200(0.60×15,000)=$4,200

     

    • High: June (15,000 hours, $13,200).

    • Low: March (13,000 hours, $12,000).


Answer (Assuming May High):
Fixed Cost = $4,400 (if recalculated with consistent high/low pairs).


answered by: anonymous